OpenStudy (veganvsmeat):

4 – 6)(3) = (4)(3) – (6)(3) A. addition property of equality B. distributive property C. multiplication property of equality D. subtraction property of equality

OpenStudy (mhchen):

a - (b)(c) = (b)(a) - (c)(a) <-- This is the distributive property, which closest matches the 4 - (6)(3) = (4)(3) - (6)(3)